A collection of dimes and nickels is worth $9.50. There are 132 coins in all.

How many of each coin are there?

dimes
nickels


Answer by AlgebraLady88(44)

We are told that a collection of dimes(x) and nickels(y) equals 132 coins in all.
The algebraic equation for that would be
x + y = 132 (1)
We are also told that all the coins add up to $9.50. The algebraic equation for that would be
0.10 x + 0.05 y = 9.50 (2)
Now we will multiply the above equation by 100 to get rid of the decimals, giving us :
10 x + 5 y = 950 (3)
Now we will multiply equation (1) by 5 , giving us
5 x + 5 y = 660 (4)
Now, we subtract : (3) - (4)
5x = 290
x = 58
This will give us y = 74.
To check , 58 ( 0.10) + 74 ( 0.05) = 9.50